# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

### v2

@@ Abstract
Recent advances in retrieval-augmented generation have significantly improved the performance of question-answering systems, particularly on factoid ‘5Ws’ questions. However, these systems still face substantial challenges when addressing ‘1H’ questions, specifically how-to questions, which are integral to decision-making processes and require dynamic, step-by-step answers. The key limitation lies in the prevalent data organization paradigm, chunk, which divides documents into fixed-size segments, and disrupts the logical coherence and connections within the context. To overcome this, in this paper, we propose Thread, a novel data organization paradigm aimed at enabling current systems to handle how-to questions more effectively. Specifically, we introduce a new knowledge granularity, termed ‘logic unit’, where documents are transformed into more structured and loosely interconnected logic units with large language models. Extensive experiments conducted across both open-domain and industrial settings demonstrate that Thread outperforms existing paradigms significantly, improving the success rate of handling how-to questions by 21% to 33%. Moreover, Thread exhibits high adaptability in processing various document formats, drastically reducing the candidate quantity in the knowledge base and minimizing the required information to one-fourth compared with chunk, optimizing both efficiency and effectiveness.
@@ 1 Introduction
Question answering (QA) is a foundational research topic in human-machine interaction (Allam & Haggag, 2012). Among the most advanced techniques, Retrieval-augmented generation (RAG) organizes external documents into fixed-size chunks and retrieves relevant knowledge to enhance QA systems (Shao et al., 2023; Trivedi et al., 2023; Jiang et al., 2023; Asai et al., 2023). This routine is particularly effective in handling the ‘5Ws’ questions, which are typically factoid requiring factual information about specific details (Yang et al., 2018; Jiang et al., 2019; Kwiatkowski et al., 2019; Stelmakh et al., 2022), such as ‘When is Shakespeare’s birthday?’. Specifically, the data organization paradigm only needs to provide chunks containing the relevant knowledge, e.g. triples or documents about the topic entity, to the RAG systems to facilitate the ‘5Ws’ questions. However, the ‘1H’ questions, derived from Aristotle’s Nicomachean Ethics (in this paper as ‘how-to’ questions, Crisp 2014), remain largely underexplored [^1] The ‘5Ws’ represent What, Why, When, Where, and Who, and the ‘1H’ stands for How. The ‘Why’ question can sometimes delve into explanations and motivations.. These questions are in high demand in practical applications such as teaching us how to write code to achieve specific goals.
Central to problem-solving (Polya & Pólya, 2014) and human learning in cognitive science (Learn, 2000), how-to questions are inherently more complex, often involve processes that require interpretation and analysis (Deng et al., 2023b). For example, correctly answering the question shown in Figure 1How to diagnose and fix a performance issue in a web application?’ involves a multi-step decision-making process, i.e., firstly checking the server load and response time, then optimizing server configuration or indexes and queries with user feedback. Such a dynamic nature necessitates the RAG systems incrementally guide users through each step, adapting to specific circumstances and providing precise and logical information. However, the prevalent data organization paradigm, chunk222‘Chunk’ refers to a general document splitting paradigm including chunks, sentences, phrases, etc. (Splitter, 2023; Chen et al., 2023; Gao et al., 2023) leaves current RAG systems (Asai et al., 2023; Shao et al., 2023) struggling to solve such questions. This chunk-based paradigm disrupts the logical flow and coherence of the content due to its inability to effectively represent the internal logic. Consequently, it returns excessive information during inference, making it difficult to maintain continuity between sequential steps, leading to fragmented responses for how-to questions and posing significant challenges for users. Overcoming these limitations necessitates a paradigm shift that accounts for the logical structure and stepwise nature of how-to questions, enabling more precise and coherent information retrieval tailored to specific needs.
@@
In this paper, we propose Thread, a new logic-based data organization paradigm designed to enhance the handling of how-to questions. The name Thread evokes the idea of ‘Pulling on the thread, the whole mystery started to unravel like a sweater.’ (Garcia & Stohl, 2011). Specifically, we introduce a new knowledge granularity named ‘logic unit’, comprising five key components and four different types (see §3.1 and §3.2). We employ a two-stage process depicted in §3.3 to extract logic units (LUs) from documents. The first, optional stage is reformulating the original documents depending on their format and style, and the second focuses on extracting and merging logic units. In this way, Thread captures connections within the documents, breaking them into more structured and loosely interconnected logic units. When answering how-to questions, the system integrated with Thread enables a dynamic interaction manner. First, it retrieves relevant LUs based on their indexed headers. Then, the body of the selected LU provides the necessary content to generate responses for the current step. With user feedback, the linker in LU dynamically connects to other LUs, allowing the system to adapt its responses until the how-to question is comprehensively addressed.
@@
We evaluate the effectiveness of Thread by conducting experiments in two open-domain, web navigation (Deng et al., 2023a) and Wikipedia Instructions (Koupaee & Wang, 2018), and one industrial setting, Incident Mitigation (Shetty et al., 2022). Experimental results demonstrate that while existing paradigms struggle with how-to questions, Thread excels at handling consecutive steps and consistently outperforms them, especially in real-world incident mitigation scenarios, with an improvement of Success Rate ranging from 21.05% to 33.33%. In addition, Thread demonstrates great efficiency and effectiveness in processing various document formats, significantly reducing the number of retrieval units to 20% and the length of tokens needed for generation to 22.60%. The main contributions of this paper include:
@@
  • We highlight the challenges faced by current RAG systems in addressing how-to questions. To address the limitation of the chunk-based paradigm, we propose Thread, a novel data organization paradigm that transforms original documents into structured, interconnected logic units.
@@
  • Integrated with Thread, our system follows a dynamic interaction mode, guiding users incrementally through each step and adapting to their specific circumstances. Our system also brings more possibilities for an automation pipeline, solving how-to questions more efficiently.
@@
  • Experimental results show that Thread significantly outperforms existing data organization paradigms across three scenarios. Additionally, Thread requires less information, while effectively handling various document formats and substantially reducing the knowledge base size.
2 Related Work
Data Organization Paradigm in RAG
The data organization process is a critical pre-stage of RAG methods where documents are processed and segmented following certain data organization paradigms. The most common data organization paradigm is splitting documents into retrieval units (Gao et al., 2023). These retrieval units vary in granularity such as phrases, sentences, propositions (Chen et al., 2023), chunks (Kamradt, 2024), etc. Coarser-grained units contain more information but introduce redundant noise, while finer-grained units have lower semantic integrity and often require retrieving more units to gather comprehensive information. However, the chunk-based data organization paradigm ignores the logical and relational connections between chunks, potentially disrupting the inherent logic flow in documents. Another paradigm constructs documents into knowledge graphs (KG), where retrieval units include entities, triplets, etc. (Gaur et al., 2022; Sen et al., 2023; He et al., 2024; Wang et al., 2024; Edge et al., 2024). These approaches emphasize semantic/lexical similarities between retrieval units, their success in factoid questions is limited when applied to how-to questions. This limitation arises because how-to questions demand logical connections between retrieval units that extend beyond mere semantic or lexical similarities.
Information Retrieved by RAG
@@ 3 Methodology
In this section, we introduce ‘logic unit’ with its components and types, then explain the process of constructing the Thread knowledge base, followed by an illustration of how Thread integrates with existing systems.
@@ 3.1 Logic Unit: Retrieval Unit of Thread
We propose a new knowledge granularity called ‘logic unit’. Different from the chunk-based paradigm, LU comprises specially designed components, especially for bridging consecutive steps when addressing how-to questions, considering the internal logic and coherence of documents.
Prerequisite. The prerequisite component acts as an information supplement, providing the necessary context to understand the LU. For example, an LU may include domain-specific terminology such as entities or abbreviations. The prerequisite explains these terms and can generate new queries to retrieve LUs with more detailed information. Without this context, passing these LUs to an LLM-based generator could lead to hallucinations. Additionally, the prerequisite can function as an LU filter, containing constraints that must be met before the LU is considered in answer generation. This filtering ensures only relevant LUs are retrieved. As shown in step LU in Figure 2, accessing the server monitor is a prerequisite for checking the server load.
@@
Body. The body contains detailed information on the LU, which is the core content fed into the LLM-based generator to generate answers. It includes specific actions or necessary information such as code blocks, detailed instructions, etc. This detailed content helps resolve the query mentioned in the header or provides a detailed explanation of the header.
Linker. The linker acts as a bridge between logic units, enabling the dynamic process of how-to questions. Unlike the chunk-based paradigm, which relies on previous retrieval units that often lack direct clues, the linker in Thread provides necessary information to generate new queries for subsequent retrieval. In Figure 2, the linker of ‘checking server’ specifies multiple possibilities after taking the action in LU body, guiding the retrieval of the next-step LU. Its format varies by LU type, serving as either a query for retrieving other LUs or an entity relationship. The edge of knowledge graph in traditional factoid questions is a special linker enabling navigation between related entities. When no further LUs are connected, the linker remains empty, isolating the current LU.
@@
Figure 2: The construction process of Thread. The left part depicts the two-stage process including reformulating documents into structured ones and extracting and merging logic units, the right part instantiates the extraction by providing an example.
@@ 3.2 Logic Unit Type
When converting documents into logic units, we distinguish how-to questions into ‘linear’ and ‘dynamic’ types, where linear how-to questions often involve a fixed sequence of steps that do not require feedback or decision points based on intermediate outcomes. Below are the common LU types identified in our experiments [^3] These LU types are summarized from our practice in experiments with industrial and public datasets. There may be additional LU types depending on the specific scenario., showing Thread is versatile enough to handle diverse types:
Step. This is the most common LU type for resolving how-to questions. Each LU body represents detailed actions, including code blocks and resolution instructions. The LU prerequisite describes the actions that need to be completed before executing the current actions. The prerequisite is especially crucial for identifying the entry point of a solution. For example, when facing a problem, there exist different ways to resolve it depending on the current situation. The prerequisite serves as a condition in the LU selection stage, filtering out LUs that do not meet the prerequisite.
Terminology. This type provides detailed explanations of domain-specific terminology. For example, terms may share the same name or abbreviation in the LU header but convey different meanings in the LU body. The prerequisites for terminology LUs describe scenarios where the terminology typically appears and linkers are usually empty unless referencing extended terminology that depends on it.
FAQ. This type provides frequently asked questions, supplementing the knowledge base. These LUs are typically isolated, with the LU body offering solutions through sequential steps that address linear how-to questions not reliant on dynamic states. They save time by avoiding the need for sequentially retrieving LUs for common questions.
Appendix. This type provides additional information relevant to the scenario of LUs, such as examples, background, lookup tables, etc. These LUs serve as supplementary knowledge for LLMs when generating responses or executable plans.
@@ 3.3 Thread: LU-based Knowledge Base
In practice, documentation is often unstructured and varies in format and style. Our approach to converting documentation into Thread involves a two-stage process to obtain LUs, as shown in the left part of Figure 2. We believe that with advancements in LLMs, this process could be streamlined into a single stage, enabling high-fidelity LU conversion with minimal hallucination.
@@
Documentation Reformulation (Optional). This stage is optional, depending on the quality of the documentation. For example, in software engineering, Troubleshooting Guides (TSGs) often vary in format, include diverse types of information, and lack readability and detail, negatively impacting productivity and service health (Shetty et al., 2022). Due to such format, where some are clearly outlined and others are disordered, we avoid directly extracting LUs from original documents. Instead, we first reformulate these documents into structured formats. Leveraging LLMs for this task, we enhance the LLMs’ in-domain understanding by providing search capabilities and domain-specific context. This is followed by a refinement step to prevent overlooking details or hallucinating information. Figure 2 (a) shows the reformulation stage and the right represents with symbols. But it is unnecessary for well-written documents like product help docs, which typically follow a linear how-to format. We illustrate the process in Table LABEL:tab:example_lu by example and provide prompts in Appendix B.
@@
LU Extraction and Merge. After reformulation, multiple LUs of varying types can be extracted from a single structured document (shown in Figure 2 (b)). Unlike chunk-based data organization commonly with fixed chunk sizes, LU granularity depends on content. For example, solutions to linear how-to questions typically form a single path from start to completion, with interconnected steps and no multiple execution outcomes encapsulated in one LU, such as an FAQ LU. However, for dynamic how-to questions with multiple possible outcomes, it is better to have one step per LU (Step LU), with Linkers navigating to the next LUs. Note that in dynamic how-to questions, not every step has multiple execution outcomes. If only one next step exists, the LUs can be merged to include both current and subsequent steps. Additionally, LUs with similar Headers and Bodies should be merged, extending the Prerequisite and Linker.
@@
LU update. In industry, documentation is often updated with each product version release. When this happens, we redo the above steps for the updated documentation, identifying LUs in Thread with their Meta Data and replacing outdated LUs.
As we extract and merge LUs, the collection of LUs from all documents forms the knowledge base, Thread. This LU-based knowledge base serves as an essential component compatible with the current RAG system and even makes it possible for automation.
3.4 Integrate Thread with QA System
To demonstrate how Thread works, we use dynamic how-to questions as an example. Figure 3 shows the RAG system incorporating our Thread data organization paradigm. The Retriever and LLM-based Generator are inherited from the original RAG system. LUs are indexed by their Headers. When an initial how-to question is submitted, the Retriever identifies the top-K most relevant LUs based on query-header similarity. The Selector then checks the prerequisites of these LUs and filters out those that do not meet the current prerequisites, derived from the initial question or any available chat history. If no current prerequisite is provided, the system can prompt the user, e.g., ‘Before doing …, have you tried …?’, to obtain the current prerequisite for LU filtering. After selection, the body of the LUs is fed into the LLM-based generator to produce an answer. If an execution engine is available, actions can be executed automatically; otherwise, the answer/action is presented to the user for manual execution. Once the action is executed, the Linker matches one of the possible outcomes and generates a new query for the next retrieval round.
Unlike traditional RAG systems, Thread-enabled systems can be potentially fully or semi-automated when integrated with execution engines. This integration offers greater automation and flexibility, as updating LUs automatically updates the system, compared to manually designed pipelines (see §3.3).
Figure 3: The RAG system integrated with Thread. It retrieves relevant LUs based on query-Header similarity and filters out LUs that do not meet the current prerequisites. The selected LUs are passed to LLMs to generate actions for execution. After execution, the Linker matches results and generates a new query for the next retrieval iteration.
4 Experimental Setup
4.1 Scenarios and Datasets
We evaluate Thread on two open-domain scenarios: Web Navigation (Mind2Web (Deng et al., 2023a)) and Wikipedia Instructions (WikiHow (Koupaee & Wang, 2018)), and one industrial setting: Incident Mitigation (IcM (Shetty et al., 2022; An et al., 2024)). We provide one example of each dataset in Appendix B to better demonstrate linear and dynamics how-to questions.
Web Navigation. Mind2Web (Deng et al., 2023a) is a dataset designed for web agents to perform complex tasks on real-world websites based on language instructions. Each task involves a ‘dynamic how-to question’, with multiple possible outcomes depending on the state of the executed actions.
Wikipedia Instructions. WikiHow [^4] https://www.wikihow.com is a platform containing numerous articles that offer step-by-step guidance on various procedural tasks. Each article is typically titled ‘How to’ and includes a brief task description, followed by a linear, fixed sequence of steps.
Incident Mitigation. Incident Mitigation (Shetty et al., 2022; An et al., 2024; Jiang et al., 2024) is essential for operating large-scale cloud services, where engineers use Troubleshooting Guides (TSGs) to address incidents. Each step in incident mitigation can yield different outcomes depending on the state, making it suitable for testing Thread on ‘dynamic how-to questions’. Unlike the other open-domain datasets, we perform a human evaluation involving twenty on-call engineers (OCEs) responsible for incident mitigation. We collect five incidents, classified into two simple and three hard ones based on their mitigation steps in history. Each OCE is tasked with mitigating all five incidents, using randomly ordered baselines per incident to avoid familiarity bias. The RAG system initiates automated mitigation for each incident; if it encounters a failure at any step, an OCE intervenes to address the issue before the system resumes automated procedures [^5] Varying across new-hire and experienced OCEs. Mitigating one incident costs each OCE around 30 minutes to 1 hour manually, and we collect five incidents to ensure consistency, engagement, and ethical treatment. Note that one OCE’s data was contaminated during the experiment, so we removed that OCE’s data..
4.2 Documents for Retrieval
Dataset #Docs #Tasks #Steps #Chunks #LUs
@@ 4.2 Documents for Retrieval
Table 1: The statistics of datasets, including the number of documents, derived chunks, LUs, etc.
As the Mind2Web dataset lacks relevant documents, we create retrieval documents tailored to the RAG system. Assuming that each website has help docs applicable across different tasks, we use the "Cross-Task" test set, selecting examples from the training set to craft informative documents for each website. And we follow Wang et al. (2023) to brainstorm different formats of documents. More details about collecting documents are shown in Appendix C.2.
For the IcM dataset, we collect 56 TSGs from an enterprise-level engineering team responsible for a large-scale cloud platform. The selected incidents in the IcM dataset can be resolved using the knowledge provided in these TSGs. We list the statistics of our datasets in Table 1.
4.3 Baselines

### v3

@@ Abstract
Recent advances in retrieval-augmented generation (RAG) have substantially improved question-answering systems, particularly for factoid ‘5Ws’ questions. However, significant challenges remain when addressing ‘1H’ questions, specifically how-to questions, which are integral for decision-making and require dynamic, step-by-step responses. The key limitation lies in the prevalent data organization paradigm, chunk, which commonly divides documents into fixed-size segments, and disrupts the logical coherence and connections within the context. To address this, we propose Thread, a novel data organization paradigm enabling systems to handle how-to questions more effectively. Specifically, we introduce a new knowledge granularity, ‘logic unit’ (LU), where large language models transform documents into more structured and loosely interconnected LUs. Extensive experiments across both open-domain and industrial settings show that Thread outperforms existing paradigms significantly, improving the success rate of handling how-to questions by 21% to 33%. Additionally, Thread demonstrates high adaptability across diverse document formats, reducing retrieval information by up to 75% compared to chunk, and also shows better generalizability to ‘5Ws’ questions, such as multi-hop questions, outperforming other paradigms.
Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation
Kaikai An1,2, Fangkai Yang2, Liqun Li2, Junting Lu1,2, Sitao Cheng2,3, Shuzheng Si1, Lu Wang2, Pu Zhao2, Lele Cao2, Qingwei Lin2, Saravan Rajmohan2, Dongmei Zhang2, Baobao Chang1 1 Peking University 2 Microsoft 3 Nanjing University ankaikai@stu.pku.edu.cn,fangkaiyang@microsoft.com,chbb@pku.edu.cn thread.github.io
@@ 1 Introduction
Question answering (QA) is a foundational research topic in human-machine interaction (Allam and Haggag, 2012). Among the most advanced techniques, retrieval-augmented generation (RAG) enhances QA systems by organizing external documents into fixed-size chunks and retrieving relevant knowledge (Shao et al., 2023; Trivedi et al., 2023; Jiang et al., 2023; Asai et al., 2023). This routine is particularly effective in handling the ‘5Ws’ questions, such as ‘When is Shakespeare’s birthday?’, which typically require the data organization paradigm providing chunks containing the relevant knowledge, e.g. triples or documents about the topic entity (Yang et al., 2018; Jiang et al., 2019; Kwiatkowski et al., 2019; Stelmakh et al., 2022). However, the ‘1H’ questions, derived from Aristotle’s Nicomachean Ethics (‘how-to’ questions, Crisp 2014), remain largely underexplored [^1] The ‘5Ws’ represent What, Why, When, Where, and Who, and the ‘1H’ stands for How. These questions are in high demand in practical applications such as teaching us how to write code to achieve specific goals.
Central to problem-solving (Polya and Pólya, 2014) and human learning (Learn, 2000), how-to questions inherently involve complex processes that require interpretation and analysis (Deng et al., 2023b). For example, answering the question in Figure 1How to diagnose and fix a performance issue in a web application?’ involves a step-by-step decision-making process, i.e., first checking server load and response time, followed by optimizing server configuration based on user feedback. This dynamic, stepwise nature necessitates RAG systems to guide users through each step, adapting to specific contexts and providing precise and logical information. However, prevalent chunk-based data organization paradigm [^2] The term ‘chunk’ here refers to a general document splitting paradigm including chunks, sentences, phrases, etc. (Splitter, 2023; Chen et al., 2023; Gao et al., 2023), which divides documents into fixed segments, disrupts the logical coherence of content. As a result, RAG systems (Asai et al., 2023; Shao et al., 2023) struggle with how-to questions, often generating excessive, fragmented information that fails to maintain continuity between steps. To address this, a paradigm shift is needed, one that preserves the logical structure and stepwise nature of how-to questions.
@@
In this paper, we propose Thread, a new logic-based data organization paradigm designed to handle how-to questions. The name Thread evokes the idea of ‘Pulling on the thread, the whole mystery started to unravel like a sweater.’ (Garcia and Stohl, 2011). Specifically, we introduce a new knowledge granularity named ‘logic unit’, comprising five key components and four different types (see §2.1 and §2.2). We employ a two-stage process (depicted in §2.3) to extract logic units (LUs) from documents. The first, optional stage is reformulating the original documents depending on their format and style, and the second focuses on extracting and merging LUs. In this way, Thread captures connections within the documents, breaking them into more structured and loosely interconnected logic units. When answering how-to questions, the system integrated with Thread enables a dynamic interaction manner. First, it retrieves relevant LUs based on their indexed headers. Then, the body of the selected LU provides the necessary content to generate responses for the current step. With user feedback, the linker in LU dynamically connects to other LUs, allowing the system to adapt its responses until the how-to question is comprehensively addressed.
@@
We evaluate the effectiveness of Thread through experiments in two open-domain, Web Navigation (Deng et al., 2023a), Wikipedia Instructions (Koupaee and Wang, 2018), and one industrial setting, Incident Mitigation (Shetty et al., 2022). Experimental results demonstrate that while existing paradigms struggle with how-to questions, Thread excels at handling consecutive steps and consistently outperforms them, particularly in real-world incident mitigation scenario, with success rate improvements ranging from 21.05% to 33.33%. Additionally, Thread shows great superiority in processing diverse document formats, reducing both the number of retrieval units and the token length required for generation. Finally, we further validate the generalizability of Thread on open-domain tasks, where it outperforms the chunk-based paradigm in handling multi-hop questions. The main contributions of this paper include:
@@
  • We highlight the challenges faced by current RAG systems in addressing how-to questions. To address the limitation of chunk-based paradigm, we propose Thread, a novel data organization paradigm that transforms original documents into structured, interconnected logic units.
@@
  • Integrated with Thread, our system follows a dynamic interaction manner, guiding users incrementally through each step and adapting to their specific circumstances. Our system also brings more possibilities for an automation pipeline, solving how-to questions more efficiently.
@@
  • Experimental results demonstrate that Thread significantly outperforms existing data organization paradigms across three scenarios. Furthermore, Thread efficiently handles various document formats, reducing the retrieval burden, and generalizes effectively to open-domain tasks.
@@ 2 Methodology
In this section, Thread leverages the internal logic and coherence of documents to construct the knowledge base, making it especially effective for addressing how-to questions. When converting documents into logic units, we classify how-to questions into two types: linear and dynamic. Linear how-to questions involve a fixed sequence of steps that do not require feedback or decision-making based on intermediate results. In contrast, dynamic how-to questions require decision-making at each step, with the process adapting dynamically based on previous actions, such as the example in Figure 1.
Below we first introduce ‘logic unit’ (LU) with its components and types, then explain how to extract LUs and construct the Thread knowledge base, followed by an illustration of how Thread integrates with existing systems.
@@ 2.1 Logic Unit: Retrieval Unit of Thread
Unlike the chunk-based paradigm, the logic unit consists of specially designed components that maintain the internal logic and coherence of a document, particularly when bridging consecutive steps in how-to questions. LUs serve as retrieval units that replace traditional chunks in RAG systems, ensuring a more structured and contextually relevant approach to information retrieval. Examples of LU in Table LABEL:tab:example_lu further illustrate such structure.
Prerequisite. The prerequisite component acts as an information supplement, providing the necessary context to understand the LU. For example, an LU may include domain-specific terminology such as entities or abbreviations. The prerequisite explains these terms and can generate new queries to retrieve LUs with more detailed information. Without this context, passing these LUs to an LLM-based generator could lead to hallucinations. Additionally, the prerequisite can function as an LU filter, containing constraints that must be met before the LU is considered in answer generation. This filtering ensures only
@@
Body. The body contains detailed information on the LU, which is the core content fed into the LLM-based generator to generate answers. It includes specific actions or necessary information such as code blocks, detailed instructions, etc. This detailed content helps resolve the query mentioned in the header or provides a detailed explanation.
Linker. The linker acts as a bridge between logic units, enabling the dynamic process of how-to questions. Unlike the chunk-based paradigm, which relies on previous retrieval units that often lack direct clues, the linker in Thread provides necessary information to generate new queries for subsequent retrieval. Its format varies by LU type, serving as either a query for retrieving other LUs (“If-Then" condition) or an entity relationship. The edge of knowledge graph in traditional factoid questions is a special linker enabling navigation between related entities. When no further LUs are connected, the linker remains empty, isolating the current LU.
@@ 2.2 Logic Unit Type
When converting documents into logic units, Thread expands its scope beyond solving how-to questions, as documents often contain more than just solutions to these types of questions. Below are the common LU types identified in our experiments [^3] These LU types are summarized from our experiments with both industrial and public datasets. Additional LU types are provided in Appendix LABEL:appendix_lu_type, and more may emerge depending on the specific scenario., demonstrating Thread’s versatility in handling a wide range of document types:
Step. This is the most common type of LU for resolving how-to questions. Each LU body encapsulates detailed actions, including code snippets and resolution instructions. The LU prerequisite specifies the conditions or actions that must be completed prior to executing the current LU, serving a critical role in determining the appropriate entry point for a solution. For instance, when encountering a problem, multiple resolution paths may exist depending on the specific context. The prerequisite functions as a filtering mechanism during the LU selection process, eliminating candidates that do not satisfy the required conditions. Samely, the linker acts as a bridge between steps, directing the flow to subsequent LUs based on the outcome of the current step.
Terminology. This type provides detailed explanations of domain-specific terminology. For example, terms may share the same name or abbreviation in LU header but convey different meanings in LU body. The prerequisites for terminology LUs describe scenarios where the terminology typically appears and linkers are usually empty unless referencing extended terminology that depends on it.
Figure 2: The upper part shows the construction process of Thread, a two-stage process including reformulating documents into structured ones (a) and extracting and merging logic units (b). The bottom (c) illustrates the RAG system integrated with Thread. It retrieves relevant LUs based on query-Header similarity and filters out LUs that do not meet the current Prerequisites. The selected LUs are passed to LLMs to generate actions based on Body for execution. After execution, the Linker matches the results and generates a new query for the next retrieval iteration.
@@ 2.3 Thread: LU-based Knowledge Base
In practice, documentation is often unstructured and varies in format and style. Our approach to converting documentation into Thread involves a two-stage process to obtain LUs, as shown in the upper part of Figure 2.
@@
Documentation Reformulation. This stage is optional, depending on the quality of the documentation. For example, in software engineering, Troubleshooting Guides (TSGs) often vary in format, include diverse types of information, and lack readability and detail, negatively impacting productivity and service health (Shetty et al., 2022). Due to such format, where some are clearly outlined and others are disordered, we avoid directly extracting LUs from original documents.
Instead, we first reformulate these documents into structured formats. By leveraging LLMs, we enhance the LLMs’ in-domain understanding by providing search capabilities and domain-specific context. This is followed by a refinement step to prevent overlooking details or hallucinating information. Figure 2 (a) shows the reformulation stage. But it is unnecessary for well-written documents like product help docs, which typically follow a linear how-to format. We instantiate this process by example in Appendix LABEL:appendix_instantiation.
@@
LU Extraction and Merge. After reformulation, multiple LUs of varying types can be extracted from a single structured document (shown in Figure 2 (b)). Unlike chunk-based data organization commonly with fixed chunk sizes, LU granularity depends on content. For example, solutions to linear how-to questions typically form a single path from start to completion, with interconnected steps and no multiple execution outcomes encapsulated in one LU, such as an FAQ LU. However, for dynamic how-to questions with multiple possible outcomes, it is better to have one step per LU (Step LU), with Linkers navigating to the next LUs. Note that in dynamic how-to questions, not every step has multiple execution outcomes. If only one next step exists, the LUs can be merged to include both current and subsequent steps. Additionally, LUs with similar Headers and Bodies should be merged, extending the Prerequisite and Linker. More details about merging are provided in Appendix LABEL:tab:lu_merge.
@@
LU update. In industry, documentation is often updated with each product version release. When this happens, we redo the above steps only for the updated documentation, identifying LUs in Thread with their Meta Data and replacing outdated LUs. As we extract and merge LUs, the collection of LUs from all documents forms the knowledge base, which serves as an essential component compatible with the current RAG system.
Dataset #Docs #Tasks #Steps #Chunks #LUs Dynamic Executable
@@
Table 1: The statistics and characteristics of datasets, including the number of documents, LUs, etc.
2.4 Integrate Thread with QA System
To demonstrate how Thread works, we use dynamic how-to questions as an example. Figure 2 (c) shows the RAG system incorporating our Thread data organization paradigm.
LUs are indexed by their Headers. When an initial how-to question is submitted, the Retriever identifies the top-K most relevant LUs based on query-header similarity. The LLM-based Selector then checks the prerequisites of these LUs and filters out those that do not meet the current prerequisites, derived from the initial question or any available chat history. After selection, the body of the LUs is fed into the LLM-based Generator to produce an answer. If an execution engine is available, actions can be executed automatically; otherwise, the answer/action is presented to the user for manual execution. Once the action is executed, the Linker matches one of the possible outcomes and generates a new query for the next retrieval round.
Unlike traditional RAG systems, Thread-enabled systems can be potentially fully or semi-automated when integrated with execution engines. This integration offers greater automation and flexibility, as updating LUs automatically updates the system, compared to manually designed pipelines.
3 Experiments
3.1 Scenarios and Datasets
We evaluate Thread on two open-domain scenarios: Web Navigation, Wikipedia Instructions, and one industrial setting: Incident Mitigation.
Web Navigation. Mind2Web (Deng et al., 2023a) is a dataset designed for web agents to perform complex tasks on real-world websites based on language instructions. Each task is treated as a multi-choice question. At each step, the input consists of HTML code, an instruction, and a set of choices, while the output is the selected choice, operation, and an optional value. We treat each one as a ‘dynamic how-to question’, with multiple possible outcomes depending on executed actions.
Wikipedia Instructions. WikiHow (Koupaee and Wang, 2018) [^4] https://www.wikihow.com provides step-by-step procedural guides, each titled ‘How to’, with a fixed sequence of steps. This dataset is used to evaluate Thread on linear how-to questions.
Incident Mitigation. IcM (Shetty et al., 2022; An et al., 2024) is critical for managing large-scale cloud services, where engineers rely on Troubleshooting Guides (TSGs) to resolve incidents. Each step can lead to different outcomes based on system states, making this dataset suitable for testing Thread on dynamic how-to questions. Unlike the open-domain datasets, we conduct a human evaluation with twenty on-call engineers (OCEs) responsible for incident mitigation. Each OCE mitigates five incidents, with baseline methods randomly assigned per incident to prevent familiarity bias. Our RAG system attempts automated mitigation, and if a failure occurs, an OCE intervenes before the system resumes [^5] Varying across new-hire and experienced OCEs. Note that one OCE’s data was contaminated during the experiment, so we removed that OCE’s data..
We collect documents from open sources, enterprises, and LLMs across scenarios. Appendix A.1 outlines the collection process, Appendix A.2 provides dataset examples, and Table 1 summarizes dataset statistics.
3.2 Baselines and Metrics